Fermi National Laboratory
 

Accelerator Update for February 6, 2006 - February 8, 2006

 

Monday February 6

The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash, and with NuMI and MiniBooNE taking beam.

Operations established store 4628 at 8:22 AM with an initial luminosity of 115.25E30.

The Meson MCenter experiment resumed taking beam at 8:40 AM. MTest came back on a little later.

Operations prepared NuMI for access at 10:57 AM.

The Antiproton Source (Pbar) resumed stacking at 12:56 PM.

At 1:16 PM, the Accumulator quadrupole (A:LQ) that had tripped off previously on a ground fault, tripped off again. By 1:41 PM, Operations had prepared Pbar for another access to investigate the problem.

At 3:38 PM, Recycler experts reported a problem with their Electron Cooling. They were accessing the Pelletron.

Pbar experts resolved the problem and began a study period around 4 PM.

Pbar resumed full intensity stacking at 5:07 PM.

Operations reported at 7:14 PM that Pbar hit a new stacking record of 17.4 mA/hr for an hour.

Operations began preparing the Recycler to stash antiprotons at 9:03 PM. The stash ended and Pbar resumed stacking at 9:51 PM.

MCenter resumed taking beam at 10:14 PM.

Tuesday February 7

The midnight shift began with Operations monitoring store 4628 and the stash, with Pbar stacking, and with NuMI, MiniBooNE, and MCenter taking beam.

The Recycler began preparing to stash at 12:59 AM. The stash ended. Pbar, MTest, and MCenter resumed taking beam at 1:36 AM.

The Recycler began preparing to stash at 4:24 AM. The stash ended. Pbar, MTest, and MCenter resumed taking beam by 5:02AM.

The Recycler began stashing around 10 AM. The stash ended and MTest and MCenter resumed taking beam. Pbar resumed stacking at 10:35 AM.

At 10:40 AM, RF experts took the Main Injector (MI) RF system down for an hour to measure and manipulate the mid-level RF.

TeV experts began an end-of-store study at 11:01 AM.

Operations terminated store 4628 at 12:24 PM.

NuMI resumed taking beam at 1:05 PM.

At 1:13 PM, MTest and MCenter lost its beam permit due to a power outage at the MS2 service building. Operations called the Duty Electrician, who had the power back on by 1:36 PM.

A Tevatron Electron Lens (TEL) expert shut off the TEL for repairs at 1:48 PM.

At 2:35 PM, a Linac RF station (KRF2) alarmed on an internal vacuum alarm. Operations contacted experts.

At 3:06 PM, Operations reported that after the MS2 power was restored, operators began turning equipment back on, but the vacuum readbacks for M01 and M02 were showing alarms. Operators contacted vacuum experts. The experts said that the alarms were not real.

Operations prepared D-Zero (D0) for access at 4:12 PM.

At 4:28 PM, NuMI tripped off due to a beam problem. Experts investigated and then accessed MI65 at 5:30 PM.

D0 came out of access at 5:40 PM.

The TEL expert reported at 5:43 PM that he replaced a faulty “HV Off” switch.

Operations put the TeV through a dry squeeze at 6 PM.

NuMI came out of access at 6:17 PM.

Operations put the TeV into shot setup at 6:30 PM.

Operations resumed sending beam to NuMI at 6:36 PM.

Operations established store 4631 at 8:48 PM with an initial luminosity of 72.63E30. The Recycler began stashing antiprotons.

The stashed ended at 9:29 PM. Pbar resumed stacking at 9:48 PM.

Wednesday February 8

The midnight shift began with Operations monitoring store 4631 and the stash, with Pbar stacking, and with MTest, MCenter, NuMI, and MiniBooNE taking beam.

A Recycler expert reported at 12:43 AM that they had fixed the Pelletron and that Electron Cooling was functional.

The Recycler began preparing to stash at 1:02 AM. The stash ended and Pbar resumed stacking at 1:38 AM. MCenter and MTest resumed taking beam.

The Recycler began preparing to stash at 4:45 AM. The stash ended and MCenter resumed taking beam at 5:26 AM. Pbar and MTest resumed taking beam.

Operations turned off NuMI beam 6:17 AM for a scheduled access.

The Plans for Wednesday

The plans for today are to stack, stash, and store.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/08/2006    email Fermilab